The number of real roots of the equation 2x4 + 5x2 + 3 = 0 is what

Answer by scott8148(6628)

factoring __ (2x^2+3)(x^2+1)=0

2x^2+3=0

x^2+1=0

the discriminants of both quadratics are negative; which means no real roots